How many gallons of a 20% antifreeze solution and how many gallons of a 50% anitfreeze solution should be mixed to get 30 gallons of a 40% antifreeze solution?
Answer by mananth(16946)

component percent ---------------- quantity
Antifreeze I 0.2 ---------------- x ml
Antifreeze II 0.5 ---------------- 30 - x ml
Mixture 0.40 ---------------- 30

0.2 x + 0.5 ( 30 - x ) = 30.00 * 0.40

0.2 x + 15 - 0.5 x = 12.00
0.2 x - 0.5 x = 12 - 15
-0.3 x = -3
/ -0.3
x = 10 gal
20.00% Antifreeze I 10 gal

50.00% Antifreeze II 20 gal